OpenStudy (anonymous):

divide as indicated x+3/3 divided by 3x+9/2

OpenStudy (lgbasallote):

\[\huge \frac{\frac{x+3}3}{\frac{3x+9}2} \implies \frac{\frac{x+3}3}{\frac{3(x+3)}2}\] does that help?

OpenStudy (anonymous):

no

OpenStudy (anonymous):

next step : \[\frac{ x+3 }{ 3} \times \frac{ 2 }{ 3(x+3) }\left[\begin{matrix}? & ? \\ ? & ?\end{matrix}\right]\]

OpenStudy (anonymous):

i really have no idea i got 7 when i divided x+3/3

OpenStudy (anonymous):



OpenStudy (anonymous):

thank u